### v1

@@ 1 Introduction
Despite the rapid success of deep neural networks in the IMDL fields [10, 43, 11], existing models suffer from inconsistent training and evaluation protocols, supported by Tables in Appendix B.1. These inconsistencies result in incompatible and unfair comparisons, yielding insufficient and misleading experimental outcomes. Hence, establishing a unified and comprehensive benchmark is the foremost concern in the IMDL field. However, establishing this benchmark is far more than straightforward protocol unification or simply model re-training. First, the training code for most state-of-the-art (SoTA) works is not publicly available, and the source code for some SoTA works is totally unreleased [26]. Second, IMDL models commonly incorporate diverse low-level features [41, 4, 13] and complex loss functions, requiring highly customized model architecture and decoupled pipeline design for efficient reproduction. Existing frameworks, like OpenMMLab [^1] https://openmmlab.com/ and Detectron2222https://github.com/facebookresearch/detectron2, heavily rely on the registry mechanism and tightly coupled pipelines. This conflict leads to severe efficiency issues while reproducing IMDL models under existing frameworks and results in monolithic model architecture with extremely high coding load and low scalability. Consequently, a comprehensive IMDL benchmark is yet to be built.
To address this issue, we introduce IMDL-BenCo, the first comprehensive IMDL benchmark and modular codebase. IMDL-BenCo: i) features a modular codebase with four components: data loader, model zoo, training script, and evaluator; the model zoo contains customizable model architecture includes SoTA models and backbone models. The loss design is also isolated within the model zoo, while other components are standardized by the interface and are highly reusable; this approach mitigates conflicts between model customization and coding efficiency;  ii) fully implements or incorporates training code for 8 SoTA IMDL models and establishes a comprehensive benchmark with 2 sets of standard training and evaluation protocols, 15 GPU-accelerated evaluation metrics, and 3 kinds of robustness evaluation, and iii) conducts in-depth analysis based on the established benchmark and codebase, offering new insights into IMDL model architecture, dataset characteristics, and evaluation standards. This benchmark and codebase represent a significant leap forward in calibrating the current progress in the IMDL field and can inspire future breakthroughs.
@@ 2 Related Works
IMDL Model Arthitectures. The key to IMDL is identifying the artifacts created by manipulation. Artifacts are considered manifest on the low-level feature space. Therefore, almost all existing IMDL models share the "backbone network + low-level feature extractor" paradigm. For example, SPAN [18] and ManTra-Net [41] use VGG [33] as the backbone of their models and combine SRM [47] and BayarConv [1] filters to obtain low-level features of the image. MVSS-Net [4] combines a Sobel [34] operator, which extracts edge information, and a BayarConv on its ResNet-50 [16] backbone to extract image noise. Detailed information about each model can be found in Table 1. Various low-level feature extractors lead to various loss functions and extremely customized model architectures. As a result, reproducing IMDL models within the existing frameworks is inefficient. The high coupling between various loss functions and training architectures also makes it extremely difficult to extend different model training frameworks. The differences between training frameworks further increase the difficulty of model reproduction. This tight coupling also severely impacts algorithm innovation and rapid iteration.
@@
IML-ViT[26] Arxiv Vision Transformer[9] Sobel Operator https://github.com/SunnyHaze/IML-ViT
Inconsisitent Training and Evaluation Protocols. Besides model reproducing difficulties, so far, there exist multiple strikingly different protocols for training and evaluating IMDL models. MVSS-Net, CAT-Net, and TruFor were pre-trained on the ImageNet [6] dataset. SPAN [18], PSCC-Net [24], CAT-Net [21], and TruFor [21] were trained using synthetic datasets. Additionally, TruFor used a large number of original images from popular photo-sharing websites Flickr and DPReview to train its Noiseprint++ Extractor. MVSS-Net [4] and IML-ViT [26] were trained on the CASIAv2 [8] dataset. on the other hand, NCL [45] did not use pre-training and was trained on the NIST16 [12] dataset. The detailed training and evaluation protocols for the models are explained in Appendix B.1. Considering the IMDL benchmark datasets are all tiny-sized (a few hundred to a few thousand images) [45], the substantial differences in training datasets make it inevitable that models using large training sets or pre-training will perform exceptionally well on other evaluation sets, posing a great challenge to the fairness of model performance evaluation. Besides, as shown in Table 1, most models do not fully open-source their code. Their results are hard to calibrate and can be highly misleading for new IMDL researchers.
@@ 3.2 Model Zoo
The model zoo currently consists of 8 SoTA models, 4 backbone models built with out-of-the-box vision backbones and 5 feature extractor modules commonly used for IMDL tasks. It is important to emphasize that we aim for all models to be trained using the same training script (standardization), while also being adaptable to all SoTA IMDL models (customization). As shown in Figure 1, we integrate the loss function computation within the forward function of the model. Through a unified interface, we pass in the required information, such as images and masks, and output the prediction results, the loss for backpropagation, and any losses, intermediate features, and images that need to be visualized. Therefore, for new IMDL methods, users only need to add the model scripts with loss design into the model zoo, and it will seamlessly integrate with all other components in IMDL-BenCo. By effectively reproducing current SoTA models using this framework, we demonstrate that we have successfully balanced the conflict between standardization and customization.
1) SoTA models. As shown in Table 1, we have faithfully reproduced 8 mainstream IMDL SoTA models, adhering to the settings from the original work. Wherever possible, we used the publicly available code, making only the necessary interface modifications. For models lacking publicly available code, we implemented them based on the settings described in their respective papers. Implementation details for each model are listed in Appendix B.2. 2) Backbone models: As classification and segmentation tasks, mainstream IMDL algorithms make extensive use of existing visual backbones, and the performance of these backbones also impacts the performance of the implemented algorithms. Therefore, we have adapted widely used vision backbones including ResNet [16], U-Net [29], ViT [9], Swin-Transformer [25], and SegFormer [42] into IMDL-BenCo as backbones. 3) Feature extractor modules: Currently, several standard feature extractors are widely used in IMDL tasks. We have implemented 5 mainstream feature extractors as $nn.module$, which include discrete cosine transform (DCT), fast Fourier transform (FFT) [2], Sobel operator [34], BayarConv [1], and SRM filter [47]—allowing seamless integration with our backbone models with registry mechanism for Managing large-scale experiments or import directly for convenient use in subsequent research.
@@ 3.4 Evaluators
Evaluation metrics are crucial for assessing the performance of IMDL models. Yet, existing methods face two key issues: 1) metrics are often unclear, with terms like optimal-F1, permute-F1, and macro-F1444https://scikit-learn.org/stable/modules/generated/sklearn.metrics.f1_score.html#sklearn.metrics.f1_score used as F1 score anonymously, and 2) most open-source codes compute metrics on the CPU, resulting in slow processing speeds.
To address these problems, we implemented GPU-accelerated evaluators in PyTorch as standard metrics. Each evaluator handles a specific metric, including Image-level (detection) F1 score, AUC (area-under-curve), accuracy, and pixel-level (localization) F1 score, AUC, accuracy, and IOU (Intersection over Union). All algorithms can automatically adapt to shape transforms in data loader conveniently. We also explicitly implemented derivative algorithms such as inverse-F1 and permute-F1 to assess their overestimated fairness, as demonstrated in Section 5.3, emphasizing the need for cautious and explicit metric declarations in future work to ensure fair and equitable comparisons.
Using 12,554 images from the CASIAv2 dataset and four NVIDIA 4090 GPUs, we tested our evaluators’ time efficiency with $nn.Identity$ as the model, which incurs negligible computation time. The results, shown in Table 2, indicate that our algorithms significantly reduce metric evaluation time, providing a faster and more reliable tool for large-scale IMDL tasks.
@@ 4 Our Benchmark
This section primarily details the benchmark testing procedures and results.
@@ 4.1 Benchmark Settings
Datasets. Our benchmark includes 8 publicly available datasets frequently used in the IMDL field: CASIA[8], Fantastic Reality[20], IMD2020[28], NIST16[12], Columbia[17], COVERAGE[40], tampered COCO[21], tampered RAISE[21].
Evaluation Metrics. Given the size imbalance between manipulated regions and authentic regions, the pixel-level F1 score is widely used to evaluate model performance. We also measure the pixel-level F1 score of the models under two different protocols. Additionally, for models that include a detection head (DH), we report the image-level F1 scores. We also report the pixel-level F1 score robustness test results for the models under Gaussian blur, Gaussian noise, and JPEG compression.
Hardware Configurations. To mimic various hardware configurations, the experiments are conducted on three distinct servers with two AMD EPYC 7542 CPUs, 128G RAM, and contain 4×NVIDIA A40 GPUs, 6×NVIDIA 3090 GPUs, and 4×NVIDIA 4090 GPUs, respectively.
Models and Hyperparameters. Our benchmark selects 8 SoTA methods in the IMDL field as the initial batch of implemented methods. The models are: Mantra-Net[41], MVSS-Net[4], CAT-Net[21], ObjectFormer[39], PSCC-Net[24], NCL-IML[45], Trufor[13], and IML-ViT[26]. The details of the corresponding tests and hyperparameters can be found in Appendix B.2.
@@ 4.2 Benchmark Protocols
Specifically, in Protocol-MVSS, we do not use authentic images during training for methods that do not have a detection head. While in Protocol-CAT, we do not make this distinction. Additionally, we use the recommended image scaling sizes specified in each paper. We also apply consistent data augmentation techniques for all methods, such as flipping, blurring, compression and simple copy-move and inpainting operations implemented using OpenCV.
@@ 4.3 Our Benchmark Results
For Protocol-MVSS, we report the F1 scores on the COVERAGE, Columbia, NIST16, CASIAv1, and IMD2020 datasets. For Protocol-CAT, we exclude IMD2020 as it is included in the training process. We use a fixed threshold of $0.5$. The pixel-level F1 scores and image-level F1 scores are presented in Table 3 and Table 4, respectively. There are some discrepancies between the performance of our reproduced models and the results reported in the original papers. Additionally, some models showed limited performance when compared under the same protocol. We also depict the robustness test under Protocol-MVSS in Figure 2. We observed significant robustness variations in some models compared to their original papers, demonstrating the necessity and fidelity of a unified framework.
@@ 5.1 Low-level Feature Extractors and Backbones
All combined models are trained on the CASIAv2 dataset for 200 epochs with an image size of 512×512. Detailed experimental settings can be found in Appendix B.5.1. They are then evaluated on four distinct datasets—CASIAv1, Columbia, NIST16, and Coverage—to assess their generalization capabilities. The average F1 score for each setting across the four datasets is reported in Table 6.
Table 6: Comparison of Generalization Performance Across Different Backbones and Feature Extractors. The Average F1 Score represents the mean F1 score across four datasets.
@@
Average F1 Score 0.312 0.078 0.094 0.089 0.075 0.097 0.227 0.144 0.162 0.207 0.142 0.226
Experiments indicate that specific feature extractors, such as BayarConv and Sobel, can negatively impact model performance. In contrast, the ResNet model, when equipped with DCT, FFT, and SRM feature extractors, shows improved performance. The ViT model, when equipped with feature extractors, tends to underfit and may require more epochs to converge. A better feature fusion method might eliminate the current issues with the ViT model. For the Swin Transformer, adding feature extractors can lead to overfitting, while the performance of the Segformer generally degrades. Detailed discussion and experiment results are detailed in Appendix B.5.2.
@@ 5.2 Dataset Bias and Cleansing Methods
Dataset Bias. Through our benchmark, we find that comparing the model performance under our protocol with the model performance after fine-tuning in their paper, on every model, a significant performance decline is observed on the NIST16 dataset. For detailed information about the NIST16 dataset and the performance decline, see Appendix B.5.3 and Table 12. However, such a huge decline does not occur on other benchmark datasets.
After thorough analysis, we find the NIST16 dataset contains "extremely similar" manipulation patterns. Then, when these extremely similar images are randomly split into the training and testing sets, models can effectively locate the manipulation regions by memorizing the extremely similar training samples. We refer to this critical dataset bias as "label leakage." Figure 3 illustrates an instance of label leakage in the NIST16 dataset.
Dataset Cleansing. Given the widespread use of NIST16 for evaluation, we propose a dataset cleansing method using the SSIM (Structural Similarity) index to ensure the cleaned NIST16-C dataset avoids overly similar images. This cleansing method is vital to address the dataset bias and calibrate the performance boost brought by label leakage. The detailed cleaning process is in the Appendix B.5.3.
@@ 5.3 Evaluation Metrics Selection
As shown in Figure 4, when the white area of the mask is large and there is a significant deviation between the model’s prediction and the mask, the inverted F1 score is much higher than F1 score. This metric affects the fairness of the evaluation.
Furthermore, using parameters such as "macro", "micro", and "weighted" when calculating F1 scores via the sklearn library is inappropriate, as it artificially inflates our F1 metrics, which is unjustified. In summary, we contend that using the F1 score with the "binary" parameter for calculation is more scientific and rigorous. We hope that future research will uniformly adopt this standard. Additionally, we discuss the current issue of AUC being overestimated in the Appendix. B.5.5.

### v2

@@ 1 Introduction
Despite the rapid success of deep neural networks in the IMDL fields [10, 45, 11], existing models suffer from inconsistent training and evaluation protocols, supported by Tables in Appendix A.1. These inconsistencies result in incompatible and unfair comparisons, yielding insufficient and misleading experimental outcomes. Hence, establishing a unified and comprehensive benchmark is the foremost concern in the IMDL field. However, constructing this benchmark is far more than straightforward protocol unification or simply model re-training. First, the training code for most state-of-the-art (SoTA) works is not publicly available, and the source code for some SoTA works is totally unreleased [27]. Second, IMDL models commonly incorporate diverse low-level features [43, 4, 13] and complex loss functions, requiring highly customized model architecture and decoupled pipeline design for efficient reproduction. Existing frameworks, like OpenMMLab [^1] https://openmmlab.com/ and Detectron2222https://github.com/facebookresearch/detectron2, heavily rely on the registry mechanism and tightly coupled pipelines. This conflict leads to severe efficiency issues while reproducing IMDL models under existing frameworks and results in monolithic model architecture with extremely high coding load and low scalability. Consequently, a comprehensive IMDL benchmark is yet to be built.
To address this issue, we introduce IMDL-BenCo, the first comprehensive IMDL benchmark and modular codebase. IMDL-BenCo: i) features a modular codebase with four components: data loader, model zoo, training script, and evaluator; the model zoo contains customizable model architecture includes SoTA models and backbone models. The loss design is also isolated within the model zoo, while other components are standardized by the interface and are highly reusable; this approach mitigates conflicts between model customization and coding efficiency;  ii) fully implements or incorporates training code for 8 SoTA IMDL models (See Table 1) and establishes a comprehensive benchmark with 2 sets of standard training and evaluation protocols, 15 GPU-accelerated evaluation metrics, and 3 kinds of robustness evaluation, and iii) conducts in-depth analysis based on the established benchmark and codebase, offering new insights into IMDL model architecture, dataset characteristics, and evaluation standards. This benchmark and codebase represent a significant leap forward in calibrating the current progress in the IMDL field and can inspire future breakthroughs.
@@ 2 Related Works
IMDL Model Architectures. The key to IMDL is identifying the artifacts created by manipulation. Artifacts are considered manifest on the low-level feature space. Therefore, almost all existing IMDL models share the "backbone network + low-level feature extractor" paradigm. For example, SPAN [18] and ManTra-Net [43] use VGG [34] as the backbone of their models and combine SRM [49] and BayarConv [1] filters to obtain low-level features of the image. MVSS-Net [4] combines a Sobel [35] operator, which extracts edge information, and a BayarConv on its ResNet-50 [16] backbone to extract image noise. Detailed information about each model can be found in Table 1. Various low-level feature extractors lead to various loss functions and extremely customized model architectures. As a result, reproducing IMDL models within the existing frameworks is inefficient. The high coupling between various loss functions and training architectures also makes it extremely difficult to extend different model training frameworks. The differences between training frameworks further increase the difficulty of model reproduction. This tight coupling also severely impacts algorithm innovation and rapid iteration.
@@
IML-ViT[27] Arxiv Vision Transformer[9] - https://github.com/SunnyHaze/IML-ViT
Inconsistent Training and Evaluation Protocols. Besides model reproducing difficulties, so far, there exist multiple strikingly different protocols for training and evaluating IMDL models. MVSS-Net, CAT-Net, and TruFor were pre-trained on the ImageNet [6] dataset. SPAN [18], PSCC-Net [25], CAT-Net [22], and TruFor [22] were trained using synthetic datasets. Additionally, TruFor used a large number of original images from popular photo-sharing websites Flickr and DPReview to train its Noiseprint++ Extractor. MVSS-Net [4] and IML-ViT [27] were trained on the CASIAv2 [8] dataset. on the other hand, NCL [47] did not use pre-training and was trained on the NIST16 [12] dataset. The detailed training and evaluation protocols for the models are explained in Appendix A.1. Considering the IMDL benchmark datasets are all tiny-sized (a few hundred to a few thousand images) [47], the substantial differences in training datasets make it inevitable that models using large training sets or pre-training will perform exceptionally well on other evaluation sets, posing a great challenge to the fairness of model performance evaluation. Besides, as shown in Table 1, most models do not fully open-source their code. Their results are hard to calibrate and can be highly misleading for new IMDL researchers.
@@ 3.2 Model Zoo
The model zoo currently consists of 8 SoTA models, 6 backbone models built with out-of-the-box vision backbones and 5 feature extractor modules commonly used for IMDL tasks. It is important to emphasize that we aim for all models to be trained using the same training script (standardization), while also being adaptable to all SoTA IMDL models (customization). As shown in Figure 1, we integrate the loss function computation within the forward function of the model. Through a unified interface, we pass in the required information, such as images and masks, and output the prediction results, the loss for backpropagation, and any losses, intermediate features, and images that need to be visualized. Therefore, for new IMDL methods, users only need to add the model scripts with loss design into the model zoo, and it will seamlessly integrate with all other components in IMDL-BenCo. By effectively reproducing current SoTA models using this framework, we demonstrate that we have successfully balanced the conflict between standardization and customization.
1) SoTA models. As shown in Table 1, we have faithfully reproduced 8 mainstream IMDL SoTA models, adhering to the settings from the original work. Wherever possible, we used the publicly available code, making only the necessary interface modifications. For models lacking publicly available code, we implemented them based on the settings described in their respective papers. Implementation details for each model are listed in Appendix A.3. 2) Backbone models: As classification and segmentation tasks, mainstream IMDL algorithms make extensive use of existing visual backbones, and the performance of these backbones also impacts the performance of the implemented algorithms. Therefore, we have adapted widely used vision backbones including ResNet [16], U-Net [30], ViT [9], Swin-Transformer [26], and SegFormer [44] into IMDL-BenCo as backbones. 3) Feature extractor modules: Currently, several standard feature extractors are widely used in IMDL tasks. We have implemented 5 mainstream feature extractors as $nn.module$, which include discrete cosine transform (DCT), fast Fourier transform (FFT) [2], Sobel operator [35], BayarConv [1], and SRM filter [49]—allowing seamless integration with our backbone models with registry mechanism for managing large-scale experiments or import directly for convenient use in subsequent research.
@@ 3.4 Evaluators
Evaluation metrics are crucial for assessing the performance of IMDL models. Yet, existing methods face two key issues: 1) metrics are often unclear, with terms like optimal-F1 [4], permute-F1 [22, 13], micro-F1 and macro-F1444https://scikit-learn.org/stable/modules/generated/sklearn.metrics.f1_score.html#sklearn.metrics.f1_score used as F1 score anonymously, and 2) most open-source codes compute metrics on the CPU, resulting in slow processing speeds.
To address these problems, we developed GPU-accelerated evaluators in PyTorch, integrated as standard metrics. Each evaluator computes a specific metric, including image-level (detection) F1 score, AUC (area-under-curve), accuracy; and pixel-level (localization) F1 score, AUC, accuracy, and IOU (Intersection over Union). All algorithms automatically adapt to shape transformations in the data loader, providing added convenience. We also explicitly implemented derived algorithms such as inverse-F1 and permute-F1 to evaluate their tendency for overestimation, as demonstrated in Section 5.3. This underscores the importance of precise and transparent metric selection in future work to ensure fair and consistent comparisons.
We experimented with 12,554 images from the CASIAv2 dataset and four NVIDIA 4090 GPUs and tested our evaluators’ time efficiency with $nn.Identity$ as the model, which incurs negligible computation time. The results, shown in Table 2, indicate that our algorithms significantly reduce metric evaluation time, providing a faster and more reliable tool for large-scale IMDL tasks.
@@ 4.1 Benchmark Settings
Datasets. Our benchmark includes eight publicly available datasets frequently used in the IMDL field: CASIA[8], Fantastic Reality[21], IMD2020[29], NIST16[12], Columbia[17], COVERAGE[42], tampered COCO[22], tampered RAISE[22]. Details of each dataset are shown in Appendix A.2.
Evaluation Metrics. Since manipulated regions are often smaller than authentic ones, the pixel-level F1 score is widely used as a suitable metric for evaluating model performance. We assess each model’s pixel-level F1 score using a fixed threshold of 0.5 across two protocols. We also evaluate all models using pixel-level AUC and IOU metrics. For models with a detection head, we additionally report image-level F1 scores. Lastly, we present robustness test results for the pixel-level F1 score under conditions of Gaussian blur, Gaussian noise, and JPEG compression.
Hardware Configurations. The experiments are conducted on three distinct servers with two AMD EPYC 7542 CPUs and 128G RAM, and contain 4$\times$×NVIDIA A40 GPUs, 6$\times$×NVIDIA 3090 GPUs, and 4$\times$×NVIDIA 4090 GPUs, respectively.
Models and Hyperparameters. Our benchmark selects eight SoTA methods in the IMDL field as the initial batch of implemented methods. The models are: Mantra-Net[43], MVSS-Net[4], CAT-Net[22], ObjectFormer[40], PSCC-Net[25], NCL-IML[47], Trufor[13], and IML-ViT[27]. The details of our minor modification, settings, and hyperparameters can be found in Appendix A.3.
@@ 4.2 Benchmark Protocols
Specifically, in Protocol-MVSS, we do not use authentic images during training for methods that do not have a detection head. While in Protocol-CAT, we do not make this distinction. Additionally, we use the recommended image scaling sizes specified in each paper. We also apply consistent data augmentation techniques for all methods, such as flipping, blurring, compression and simple copy-move and inpainting operations implemented using OpenCV [^5] https://opencv.org/.
@@ 4.3 Our Benchmark Results
For Protocol-MVSS, we report the F1 scores on the COVERAGE, Columbia, NIST16, CASIAv1, and IMD2020 datasets. For Protocol-CAT, we exclude IMD2020 as it is included in the training process. The pixel-level F1 scores and image-level F1 scores are presented in Table 3 and Table 4, respectively. The pixel-level AUC and IoU can be found in Appendix A.4. There are some discrepancies between the performance of our reproduced models and the results reported in the original papers. Additionally, some models showed limited performance when compared under the same protocol. We also depict the robustness test under Protocol-MVSS in Figure 2. We observed significant robustness variations in some models compared to their original papers, demonstrating the necessity and fidelity of a unified framework.
@@ 5.1 Low-level Feature Extractors and Backbones
All combined models are trained on the CASIAv2 dataset for 200 epochs with an image size of 512×512. Detailed experimental settings can be found in Appendix A.7.1. They are then evaluated on four distinct datasets—CASIAv1, Columbia, NIST16, Coverage, and IMD2020—to assess their generalization capabilities. The average F1 score for each setting across the four datasets is reported in Table 6.
Table 6: Comparison of Generalization Performance Across Different Backbones and Feature Extractors. The Average F1 Score represents the mean F1 score across five datasets.
@@
Average F1 Score 0.295 0.068 0.228 0.241 0.042 0.254 0.213 0.133 0.151 0.188 0.128 0.216
Experiments indicate that specific feature extractors, such as BayarConv and Sobel, can negatively impact model performance. In contrast, the ResNet model, when equipped with DCT, FFT, and SRM feature extractors, shows improved performance. The ViT model, when equipped with feature extractors, tends to underfit and may require more epochs to converge. A better feature fusion method might eliminate the current issues with the ViT model. For the Swin Transformer, adding feature extractors can lead to overfitting, while the performance of the Segformer generally degrades. Detailed discussion and experiment results are detailed in Appendix A.7.2.
@@ 5.2 Dataset Bias and Cleansing Methods
Dataset Bias. Through our benchmark, we find that comparing the model performance under our protocol with the model performance after fine-tuning in their paper, on every model, a significant performance decline is observed on the NIST16 dataset. However, such a huge decline does not occur on other benchmark datasets. After thorough analysis, we find the NIST16 dataset contains "extremely similar" manipulation patterns. Then, when these extremely similar images are randomly split into the training and testing sets, models can effectively locate the manipulation regions by memorizing the extremely similar training samples. We refer to this critical dataset bias as "label leakage." Figure 3 illustrates an instance of label leakage in the NIST16 dataset.
Dataset Cleansing. To enhance the reliability of NIST16 for evaluation, we introduce a new dataset, NIST16-C [^6] NIST16-C is available here: https://github.com/DSLJDI/NIST16-data-set-deduplication, created by applying a filtering method based on Structural Similarity (SSIM) [41]. This process helps eliminate overly similar images, reduce dataset bias, and prevent performance overestimation caused by label leakage. Further details for our analysis and the cleansing procedure can be found in Appendix A.7.3.
@@ 5.3 Evaluation Metrics Selection
Furthermore, using parameters such as "macro", "micro", and "weighted" when calculating F1 scores via the sklearn library is inappropriate, as it artificially inflates our F1 metrics, which is unjustified. We further analyze these misleading F1 metrics in Appendix A.7.4. Mixing these F1 scores anonymously would lead to significant fairness issues. In summary, we contend that using the F1 score with the "binary" parameter for calculation is more scientific and rigorous. We hope that future research will uniformly adopt this standard. Additionally, we discuss the current issue of AUC being overestimated in the Appendix A.7.5.